40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 114 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  11598   Tue Sep 15 15:01:23 2015 ranaSummaryLSCdisabling the LSC AA filters + mod to whitening

While investigating the BIO situation with the LSC machine and the iscaux2 processor last night, we wondered if maybe the Anti-Aliasing filters were mistakenly disabled. But why do we need these anyway?

Our ADCs digitize at 64 kHz and there is a digital lowpass in the IOP at 5 kHz before we downsample to 16 kHz. So mainly we're trying to prevent some aliasing at the 64 kHz IOP rate. But our analog AA filter is a 8th order ELP at 7570 Hz, so its overkill.

So, I propose that we bypas the AA via hardwiring the board and implement a 10 kHz pole in the whitening board (D990694) before the whitening by turning R127, etc. into a 0.1 uF cap. Along with the 100 Ohm series resistor, this will make a pole at ~15 kHz. Probably ought to check that the input resistor is metal film. Also, if we replace C158/C159, etc. with a 0.47 nF cap, we'll get 2 poles at 35 kHz to limit the higher frequencies from saturating.

  11597   Tue Sep 15 01:14:10 2015 ranaSummaryLSCneed to check LSC Whitening switch logic ... again

Tonight we noticed that the REFL_DC signal has gone bipolar, even though the whitening gain is 0 dB and the whitening filter is requested to be OFF.

We should check out the switch operation of several ofthe LSC channels in the daytime - where is the procedure for this diagnostic posted?

  11596   Mon Sep 14 23:12:49 2015 ericqUpdateLSC55MHz modulation phase effect on PRMI

With the adjustable delay line box installed in the 55MHz modulation path, I've measured the PRMI sensing matrix as a function of delay / relative phase between the 11MHz and 55MHz modulations. The relative frequency difference of 44MHz tells us that this should be cyclical after ~23nsec of delay, but losses in the delay cable change this; see Koji's elogs about the modulation cancellation setup for details. 

TL;DR: Nothing really changes, other than REFL33 optical gain. MICH/PRCL angles remain degenerate.


The results aren't so surprising. The demod angles for the 55MHz diodes don't even change, since the same 55MHz signal is used for the modulator and demodulators, so delaying it before the split should go unnoticed. Most of these measurements were made during the same lock stretch, PRCL on REFL11 I and MICH on AS55Q.

The only signals we would expect to change much are ones that have significant contriubtions from field products influenced by both modulations. None of the 1F PDs are like this, nor is REFL165. REFL33 is the odd man out, where the +44MHz field produced as a -11MHz sideband on the +55MHz sideband beats with the +11MHz sideband (and the same with the signs flipped). I made a simulation for the 40m poster at the March 2015 LVC meeting, but I don't think it ever made it to the ELOG. 

So:

Here are the results for the 0ns and 4ns cases, as an illustration of what changes (REFL33), and what doesn't (everything else). Again, these are calibrated to Volts out of the analog demod boards per meter of DoF motion. 

 

So, since REFL33 is the only one really changing, let's just look at it by itself:

Qualitatively, the change in magnitude looks similar to the simulation result. The demod angles fall by some roughly linear amount. The angle difference is even more stationary than predicted there, though. 

Attachment 1: PRMI_CAR_0ns.pdf
PRMI_CAR_0ns.pdf
Attachment 2: PRMI_CAR_4ns.pdf
PRMI_CAR_4ns.pdf
Attachment 3: delaySweep_nominal.pdf
delaySweep_nominal.pdf
Attachment 4: 55delay_PRMI_REFL33.pdf
55delay_PRMI_REFL33.pdf
  11595   Mon Sep 14 21:42:00 2015 ranaUpdateIOOMC Wiener + Summary

I turned on the MCF FF in the OAF today (we need to fix the labeling of the 'ON' buttons on the RHS of the screen). The performance is still good; before / after attached.

Not only is the 1 Hz performance in the MC still good, but the X & Y arm noise reduction is ~1 order of magnitude. Good to know that the filters aren't changing much with time.

Can we just leave this on all the time now? Seems to be OK and there's no visible increase in the arm noise with this on.

Also did some updates to the summary pages and added a CDS FEC tab for CPU times.

Please take a look at the summary pages and bring a list of demands to the Wednesday meeting.

Attachment 1: mcf.pdf
mcf.pdf
  11594   Mon Sep 14 16:50:12 2015 ericqUpdateLSCQuick note

Just a heads up while I'm out for a bit: the delay line is currently installed in the 55MHz modulation path. 

I'll be back later, and will revert the setup.

  11593   Mon Sep 14 10:41:03 2015 SteveUpdateVACTP2 dry pump replaced

TP2 dry fore pump sn PLE10082 was replaced at pressure 717 mTorr,  TP2 50K rpm 0.33A @ 112,677 hrs

Top seal life was 8,160 hrs

Model  SH110, Sn LP1007L556 was installed. It's fore line pressure after 30 minutes of running 38 mTorr, TP2 turbo at 50K rpm 0.18A

 

 

Attachment 1: TP2_dry_pump_replaced.png
TP2_dry_pump_replaced.png
  11592   Sun Sep 13 13:26:00 2015 ranaUpdateIOOLast Wiener MCL subtractions

When making the Wiener filter OFF/ON comparisons, we want to use the median PSD estimates, not the mean (which is what pwelch gives you).

cf. Sujan's note and Evan's follow-up

The median will be less sensitive to the transients / gltiches and will show more improvement I think.

  11591   Fri Sep 11 10:56:47 2015 SteveUpdateendtable upgradeETMY optical table feedthrough

The ETMY enclosure feedthrough - north is installed. The sealing material is hard to work with.

The upper empty blocks will be replaced by something soft to make changing cables easy.

 

Attachment 1: ETMY-Nfeedt.jpg
ETMY-Nfeedt.jpg
  11590   Thu Sep 10 09:37:34 2015 IgnacioSummaryIOOFilters left on MCL static module

The following MCL filters were left loaded in the T240-X and T240-Y FF filter modules (filters go in pairs, both on):

FM7: SISO filters for MCL elog:11541

FM8: MISO v1 elog:11547

FM9: MISO v1.1 Small improvement over MISO v1

FM10 MISO v2 elog:11563

FM5 MISO v3.1 elog:11584 (best one)

FM6 MISO 3.1.1 elog:11584 (second best one)

 

  11589   Thu Sep 10 04:23:00 2015 ericqUpdateLSCMoved LSC sensing matrix notch frequencies

Frequencies are:

  • CARM: 309.21 Hz
  • DARM: 307.88 Hz
  • MICH: 311.1 Hz
  • PRCL: 313.31 Hz
  • SRCL: 315.17 Hz

POP110 and POP22 demod angles were adjusted for DRMI lock. 

Last week, I never achieved a fully 1F lock, REFL165 was used for SRCL. Tonight, we created input matrix settings for pure 1F locking, and did some signal mixing to reduce the PRCL to SRCL coupling. The PRCL to MICH coupling was already low, since AS55 is fairly insensitive to PRCL. 

Similarly, for the 3F signals, some signal mixing of REFL33I and REFL165Q was used to reduce the PRCL to MICH coupling. The PRCL to SRCL coupling in REFL165 isn't too bad, so no compensation was done. Interestingly, in this setting, the 3F MICH and SRCL signals agree with the 1F signals on their zero crossing, so no offsets are needed. REFL33 I does need an offset, however, to match the REFL11I PRCL zero crossing. 

The DRMI acquires faster with SRCL set to 165I. Once acquired, the 1F/3F can be made smoothly, and both settings are very stable. The sensing matrix in each setting is consistent with each other. (The PRCL and SRCL lines in AS55 change, but really I shouldn't even plot them, since they're not very coherent). 

For some reason, these show a sign flip relative to last week's measurements. The relative angles are consistent, though. 

Next up is finding the right coefficient for the SRM in the MICH output matrix, when actuating on the BS. 

Attachment 1: DRMI_1F.pdf
DRMI_1F.pdf
Attachment 2: DRMI_3F.pdf
DRMI_3F.pdf
  11588   Thu Sep 10 01:09:20 2015 ranaUpdateLSCMoved LSC sensing matrix notch frequencies

We looked at the DRMI noise spectrum and chose new excitation frequencies such that the lines are lower in frequency than before (~300 Hz instead of 800 Hz) and also not in some noisy region.

New filters is saved and loaded for all LSC DOFs.

Attachment 1: NewLSCnotches.png
NewLSCnotches.png
  11587   Wed Sep 9 15:45:11 2015 ericqUpdatePEMInverted STS filters in C1OAF

Our online subtraction filters for PRC angle and MC length were trained on the raw ADC signals, so I've inverted the filters that Rana installed in the PEM filter banks in the OAF signal conditioning filter banks (C1:OAF-WIT_STS1X, etc.)

It's not perfect, since the inversion would be unstable, and thus needs a low pass. I used an ELP at 800Hz. The error in the inversion is then something like half a degree at 5Hz, which is the highest frequency we really ever subtract at. It should be ok.

  11585   Wed Sep 9 11:33:58 2015 ranaUpdateSummary PagesSummary Page updates
  • Made most plots in IOO tab only plot when MC_TRANS > 10000 using Eve's MC_LOCK state definition.
  • added the 0.03 - 0.1 Hz and 10-30 Hz bands to the PEM SEIS BLRMS tab and set the y-scales to the same as SeismicRainbowSTS.stp
  • set state PMC_LOCK in PSL tab and made some of those plots only plot when PMC trans > 0.6.
  • SUS-OL page showed me that the ETM yaw spectrum was wacky, which lead me to find that it was completely uncentered. Stop leaving the room lights ON Steve!!  angry I also set the quadrant offsets by blocking the QPD with a piece of metal (teflon doesn't work).
  • set c1summary to only plot some when X or Y arms are locked
  11584   Wed Sep 9 11:00:49 2015 IgnacioUpdateIOOLast Wiener MCL subtractions

On Thursday night (sorry for the late elog) I decided to give the MCL FF one more try. 

I first remeasured the actuator transfer function because previous measurements had poor coherence ~0.5 - 0.7  at 3 Hz. I did a sine swept to measure the TF. 

Raw transfer function:

The data is attached here: TF.zip

Then I made Wiener filters by fitting the transfer function data with coherence > 0.95 (on the left). Fitting all the data (on the right). Here are the filters:

 

The offline subtractions (high coh fit on left, all data fit on right). Notice the better IIR performance when all the TF data was fitted.

 

The online results: (these were aquired by taking five DTT measurements with 15 averages each and then taking the mean of these measurements)

 

And the subtraction performance:

 

Attachment 3: TF.zip
  11583   Tue Sep 8 20:30:44 2015 ranaUpdateIOOMC WFS relief re-commissioned

I converted our MC WFS relief from CSH to BASH today. I also added 'wait' commands and 'echo' commands so that all DoFs run in parallel nicely. It can be accessed from the MC WFS screen.

I increased the overall MC WFS gain input slider from 0.02 to 0.1 (its in the mcwfson script). The MC Trans loops now have a time constant of ~30 seconds. The relief script relieves ~90% of the MC WFS control signals in the 2 minutes that its allowed to run.

On the next upgrade, we should make it python and have it kill the relief process if the MC loses lock before relief is applied via the alignment sliders.

Attachment 1: WFSrelief.png
WFSrelief.png
  11582   Mon Sep 7 19:46:46 2015 ranaUpdatePEMSeismic BLRMS filters

As it turned out, the "STS" BLRMS filters were all a mess, so I fixed them up today:

  • BP and LP filters were non-existent for the 2 low frequency bands: 0.01-0.03 & 0.03-0.1 Hz. The 0.01-0.03 is just seeing tilt noise (its big in X & Y, but not in Z), but the other band is able to cleanly see the primary microseism at 0.06 Hz.
  • There was some mixup and some BP filter banks had low pass filters while one of the LP banks had a BP filter.
  • There were different filters between the X, Y & Z directions.
  • The low pass filters had enough ringing in the impulse response that their outputs could sometimes go negative and make the SQRT block output NaN.

After tuning:

  • All bandpass filters are 4th order Butterworth bandpass with the corners at the band edges (e.g. 1- 3  Hz)
  • All low pass are the same, just scaled by the frequency band. They have a pair of real poles and a pair of 35 deg poles. The pole frequencies are set so that there is 40 dB of attenuation at twice the frequency of the low end of the bandpass. i.e. for the 1-3 Hz band, the low pass has > 40 dB atten at 2 Hz.
  • The 3-10 and 10-30 Hz bands use the same low pass as the 1-3 Hz band, since I don't want to see aliasing in the EPICS readouts. I don't think we need faster than 1 Hz readback of the RMS.
  • Confirmed with FOTON that the impulse response for the LP filters are positive for all t >0.

The "C1:PEM-SEIS_STS_1" filter banks are currently empty, so the signal is just in ADC counts. However, by amazing luck, this seems to be the right gain (within a factor of 2) to put the signal into units of microns / second. According the the schematic (D1000749), the default gain of 110 can be switched to make the whole box just have a gain of 2 (differential in, differential out). I wonder if anyone, like Jenne, knows if this is what we have? There's no elog I found about setting the gain switch.

According to the manual, the gain is ~1175 V/(m/s). Our ADC gain should be (2^16)/(40). So:

cal_gain = 1175 * 2 * 65536 / 40  ==>> 0.26 (m/s)/counts

I have put this into the STS_1_X,Y,Z filter modules in c1pem so that these channels are now calibrated. I also put the first few s-domain poles/zeros into the filter based on the manual so that the magnitude in the 10-30 Hz band is correct-ish now.

* Does anyone know how to center the masses on this thing?

Attachment 1: T240_150907.png
T240_150907.png
  11581   Mon Sep 7 18:25:16 2015 ranaConfigurationIOOAOM stage is ready

The new stage missed the right height by ~2 mm. sad

Even if I completely bottom out the (New Focus 9071) 4-axis stage, its not short enough. So I removed the AOM from the beam and re-aligned into the PMC.

Steve, please get the aluminum piece remachined to go down by 2.5 mm so we can have some height adjustment room.

Quote:

New stage can cheeky hold the correct polarization.

Also, the turning mirror mount just after the EOM and before the AOM is a U-100 and we want it to be a Suprema for stability - let's not forget to swap that after Steve gets the mount fixed.

  11580   Mon Sep 7 16:30:56 2015 ranaHowToComputer Scripts / Programsincrease of window border size on Rossa

Frustrated by the single pixel width of the windows and how hard that makes it to drag things around, I explored StackExchange:

which showed how there is a .xml file which can be edited to increase this. I've changed the border size to 4 pixels on Rossa - its nice.devil

  11579   Fri Sep 4 20:42:14 2015 gautam, ranaUpdateCDSCheckout of the Wenzel dividers

Some years ago I bought some dividers from Wenzel. For each arm, we have x256 and a x64 divider. Wired in series, that means we can divide each IR beat by 2^14.

The highest frequency we can read in our digital system is ~8100 Hz. This corresponds to an RF frequency of ~132 MHz which as much as the BBPD could go, but less than the fiber PDs.

Today we checked them out:

  1. They run on +15V power.
  2. For low RF frequencies (< 40 MHz) the signal level can be as low as -25 dBm.
  3. For frequencies up to 130 MHz, the signal should be > 0 dBm.
  4. In all cases, we get a square wave going from 0 ~ 2.5 V, so the limiter inside keeps the output amplitude roughly fixed at a high level.
  5. When the RF amplitude goes below the minimum, the output gets shaky and eventually drops to 0 V.

Since this seems promising, we're going to make a box on Monday to package both of these. There will one SMA input and output per channel.

Each channel will have a an amplifier since this need not be a low noise channel. The ZKL-1R5 seems like a good choice to me. G=40 dB and +15 dBm output.

Then Gautam will make a frequency counter module in the RCG which can do counting with square waves and not care about the wiggles in the waveform.

I think this ought to do the trick for our Coarse frequency discriminator. Then our Delay Box ought to be able to have a few MHz range and do all of the Fast ALS Carm that we need.

Attachment 1: TEK00000.PNG
TEK00000.PNG
Attachment 2: TEK00001.PNG
TEK00001.PNG
Attachment 3: TEK00002.PNG
TEK00002.PNG
  11578   Fri Sep 4 20:06:23 2015 ranaUpdateLSCDRMI locked on 1F and 3F

Nice going. I think the LLO / LHO scheme is to acquire on 1F and then cdsutils avg to get the 3F offsets. The thinking is that that 1F signals have less intrinsic offset than the 3F signals, so we want to be use digital offsets for the 3F locks.

  11577   Fri Sep 4 15:20:31 2015 ericqUpdateLSCVertex Sensing

I've now made a collection of sensing matrix measurements. 

In all of the plots below, the radial scale is logarithmic, each grid line is a factor of 10. The units of the radial direction are calibrated into demod board output Volts per meter. The same radial scale is used on all plots and subplots.

I did two PRMI measurements: with MICH locked and excited with either the ITMS or the BS + PRM compensation. This tells us if our PRM compensation is working; I think it is indeed ok. I though I remembered that we came up with a number for the SRM compensation, but I haven't been able to find it yet. 

The CARM sensing int he PRFPMI measurement has the loop gain at the excitation frequency undone. All excitations were simultaneously notched out of all control filters, via the NotchSensMat filters. 

The angular scale is set to the analog I and Q signals; the dotted lines show the digitial phase rotation angle used at the time of measurement. 

Attachment 1: PRFMI_ITM.pdf
PRFMI_ITM.pdf
Attachment 2: PRFMI_BS.pdf
PRFMI_BS.pdf
Attachment 3: DRMI.pdf
DRMI.pdf
Attachment 4: PRFPMI.pdf
PRFPMI.pdf
  11576   Fri Sep 4 10:25:19 2015 SteveConfigurationIOOAOM stage is ready

New stage can hold the correct polarization.

DRAWING CORRECTION:  Post block height was lowered to be 1.88" from 2.0"

Attachment 1: stage_AOM-EOM.jpg
stage_AOM-EOM.jpg
Attachment 2: A-EOMholder.pdf
A-EOMholder.pdf A-EOMholder.pdf A-EOMholder.pdf A-EOMholder.pdf
  11575   Fri Sep 4 09:36:48 2015 SteveUpdateLSCIR beatnote confusion.....

 

Quote:

I was thinking that the "FOSC" product line (which is called a "coupler" instead of a "splitter/combiner") was what we wanted. 

Koji brought to my attention that the 90/10 splitters we already have are of this line. So, I rigged a few up to shine a hopefully beating pair of fields on the fiber coupled thorlabs PD. 

I was able to get ~80uW each of PSL and AUX X light on the PD, which produced a -10dBm beatnoteyes Thus, I think these FOSC splitters are indeed what we want. 

I then threw this IR beatnote at our ALS signal chain. The beatnote was too big to throw through our ~+27dB RF amps, so I just sent the -10dBm over to the LSC rack.

The IR beat spectrum is somwhat noisier from 10-100Hz, but, more interesting, is that the sub-4Hz noise is identical in the two beats, and very coherent. This excludes ALS noise arising from anything happening in the green beat optics on the PSL table.

Obviously, the high frequency noise is largely the same and coherent too, but also coherent with the AUX X PDH control signal, so it is understood. 

FOSC-2-64-50-L-1-H64F-2
Single mode coupler, 2x2, 1064nm +/-20nm, 50/50 ratio, 900micron loose tube jacket, Hi1060flex fiber, 1m fiber length, FC/APC connectors

Four of these items ordered yesterday from http://afwtechnologies.com.au/sm_coupler.html

  11574   Fri Sep 4 09:23:32 2015 IgnacioUpdateCDSModified Pentek schematic

Attached is the modifed Pentek whitening board schematic. It includes the yet to be installed 1nF capacitors  and comments. 

Attachment 1: schematic.pdf
schematic.pdf schematic.pdf schematic.pdf
  11573   Fri Sep 4 08:00:49 2015 IgnacioUpdateCDSRC low pass circuit (1s stage) of Pentek board

Here is the transfer function and cutoff frequency (pole) of the first stage low pass circuit of the Pentek whitening board.

Circuit:

R1 = R2 = 49.9 Ohm, R3 = 50 kOhm, C = 0.01uF. Given a differential voltage of 30 volts, the voltage across the 50k resistor should be 29.93 volts.

Transfer Function: 

Given by, 

H(s) = \frac{1.002\text{e}06}{s+1.002\text{e}06}

So low pass RC filter with one pole at 1 MHz.

I have updated the schematic, up to the changes mentioned by Rana plus some notes, see the DCC link here: [PLACEHOLDER]

I should have done this by hand...crying

Attachment 1: circuit.pdf
circuit.pdf
  11572   Fri Sep 4 04:12:05 2015 ericqUpdateComputer Scripts / ProgramsMATLAB down on all workstations

There seems to be something funny going on with MATLAB's license authentication on the control room workstations. Earlier today, I was able to start MATLAB on pianosa, but now attempting to run /cvs/cds/caltech/apps/linux64/matlab/bin/matlab -desktop results in the message: 

License checkout failed.
License Manager Error -15
MATLAB is unable to connect to the license server. 
Check that the license manager has been started, and that the MATLAB client machine can communicate
with the license server.

Troubleshoot this issue by visiting: 
http://www.mathworks.com/support/lme/R2013a/15

Diagnostic Information:
Feature: MATLAB 
License path: /home/controls/.matlab/R2013a_licenses:/cvs/cds/caltech/apps/linux64/matlab/licenses/license.dat:/cv
s/cds/caltech/apps/linux64/matlab/licenses/network.lic 
Licensing error: -15,570. System Error: 115

  11571   Fri Sep 4 04:05:51 2015 ericqUpdateLSCDRMI locked on 1F and 3F

Thanks to some expertly timed coffee from Ignacio, I have been able to achieve indefnite locks of the DRMI, first on a 1F/3F mix (P:REFL11, S: REFL165, M:AS55), and then purely on 3F (P:REFL33, S:REFL165, S:REFL165). MICH is currently actuated on the ITMs. 

I saved a snapshot of the current settings so I don't lose my settings. I think one thing that prevented earlier recipies from working is that whitening gains may have changed, which we don't typically note down when reporting input matrix settings

My current settings for 3F locking:

REFL33:

+30dB whitening gain, +136 demod phase

PRCL = 9 x I - 200 counts

REFL165:

+24dB whitening gain, +3 demod phase

SRCL = 1 x I, MICH = 5 x Q - 1000counts

MICH: G=-0.03; Acq FM4/5; Trig 2/3/6/9

PRCL: G=-0.003; Acq FM4/5; Trig 1/2/6/9

SRCL: G=0.2; Acq FM4/5; Trig 2/3/6/9

I've injected excitations into the control filter outputs via the LSC-FFC FMS (and notched the frequencies in the control filters themselves), and noted GPS times for offline sensing analysis. (Namely the 10 minutes following 1125398900)

Handing off to pure 3F was a little finicky at first, I needed to use some pretty large offsets in the MICH_B and PRCL_B FMs. (-1000 and -200 counts respectively). Once these offsets were found, the DRMI can acquire on 3F. Alignment is pretty important, too.  Acquiring is much faster when the loop gains are "too high." i.e. I see a fair amount of gain peaking at ~300Hz. Nevertheless, things are stable enough as is that I didn't feel like digging into reducing the gains to quieter values. 

  11570   Fri Sep 4 00:58:29 2015 ranaUpdateCDSsoldering the Generic Pentek interface board

Q and Ignacio were taking a second look at the Pentek interface board which we're using to acquire the POP QPD, ALS trans, and MCF/MCL channels. It has a differential intput, two jumper able whitening stages inside and some low pass filtering.

I noticed that each channel has a 1.5 kHz pole associate with each 150:15 whitening stage. It also has 2 2nd order Butterworth low pass at 800 Hz. Also there's a RF filter on the front end. We don't need all that low passing, so I started modifying the filters. Tonight I moved the 800 Hz poles to 8000 Hz. Tomorrow we'll move the others if Steve can find us enough (> 16) 1 nF SMD caps (1206 NPO).

After this those signals ought to have less phase lag and more signal above 1 kHz. Since the ADC is running at 64 kHz, we don't need any analog filtering below 8 kHz.

  11569   Thu Sep 3 19:52:24 2015 ranaSummarySUSSUS drift monitor

Since Andrey's SUS Drift mon screen back in 2007, we've had several versions which used different schemes and programming languages. Diego made an update back in January.

Today I added his stuff to the SVN since it was lost in the NFS disks somewhere. Its in SUS/DRIFT_MON/.

Since we've been updating our userapps directory recently to pull in the screens and scripts from the sites, we also got a copy of the Thomas Abbott drift mon stuff which is better (Diego actually removed the yellow/red functionality as part of the 'upgrade'), but more complicated. For now we have the old one. I updated the good values with all optics roughly aligned just a few minutes ago.
 

Attachment 1: 07.png
07.png
  11568   Thu Sep 3 17:15:26 2015 ericqUpdateLSCIR beatnote confusion

I was thinking that the "FOSC" product line (which is called a "coupler" instead of a "splitter/combiner") was what we wanted. 

Koji brought to my attention that the 90/10 splitters we already have are of this line. So, I rigged a few up to shine a hopefully beating pair of fields on the fiber coupled thorlabs PD. 

I was able to get ~80uW each of PSL and AUX X light on the PD, which produced a -10dBm beatnoteyes Thus, I think these FOSC splitters are indeed what we want. 

I then threw this IR beatnote at our ALS signal chain. The beatnote was too big to throw through our ~+27dB RF amps, so I just sent the -10dBm over to the LSC rack.

The IR beat spectrum is somwhat noisier from 10-100Hz, but, more interesting, is that the sub-4Hz noise is identical in the two beats, and very coherent. This excludes ALS noise arising from anything happening in the green beat optics on the PSL table.

Obviously, the high frequency noise is largely the same and coherent too, but also coherent with the AUX X PDH control signal, so it is understood. 

Attachment 1: GREENvIRbeats.pdf
GREENvIRbeats.pdf
  11567   Thu Sep 3 13:25:40 2015 ranaUpdateCDSSimulink Webview updated

added the cron script for this to megatron to run at 8:44 AM each morning. Here's the new MegaCron attached :-()-

** it takes ~13 minutes to complete on megatron

Attachment 1: crontab_150903.rtf
MAILTO=ericq@caltech.edu

# m h  dom mon dow   command
#0 */1 * * * bash /home/controls/public_html/summary/bin/c1_summary_page.sh > /dev/null 2>&1
#15 5 * * * /ligo/apps/nds2/nds2-megatron/test-restart

# MEDM Screen caps for the webpage
2,13,25,37,49 * * * * /cvs/cds/project/statScreen/bin/cronjob.sh

# op340m transplants -ericq
... 18 more lines ...
  11566   Thu Sep 3 08:53:55 2015 SteveUpdateSUSETMY Oplev laser power is falling

Rana, Steve

We investigated the ETMY oplev table set up and did not find a red herring.

Two 2 years  vs one day plot below.

ps: thanks  Q for fixing  DTT, the auto scaling is not working at sampling rate 10 min and 1 hr period?? surprise

 

Attachment 1: ETMYoplev2y.png
ETMYoplev2y.png
  11565   Thu Sep 3 02:30:46 2015 ranaUpdateCDSc1cal time reduced by deleting LSC sensing matrix

I experimented with removing somethings here and there to reduce the c1cal runtime. Eventually I deleted the LSC Sensing Matrix from it.

  • Ever since the upgrade, the c1cal has gone from 60 to 68 usec run time, so its constantly over.
  • Back when Jenne set it up back in Oct 2013, it was running at 39 usec.
  • The purely CAL stuff had some wacko impossible filters in it: please don't try to invert the AA filters making a filter with multiple zeros in it Masayuki.
  • I removed the weird / impossible / unstable filters.
  • I'm guessing that the sensing matrix code had some hand-rolled C-code blocks which are just not very speedy, so we need to rethink how to do the lockin / oscillator stuff so that it doesn't overload the CPU. I bet its somewhere in the weird way the I/Q signals were untangled. My suggestion is to change this stuff to use the standard CDS lockin modules and just record the I/Q stuff. We don't need to try to make magnitude and phase in the front end.

After removing sensing matrix, the run time is now down to 6 usec.

  11564   Thu Sep 3 02:12:08 2015 ranaUpdateCDSSimulink Webview updated

Back in 2011, JoeB wrote some entries on how to automatically update the Simulink webview stuff.

Somehow, the cron broke down over the years. I reran the matlab file by hand today and it worked fine, so now you can see the up to date models using the internet.

https://nodus.ligo.caltech.edu:30889/FE/

  11563   Thu Sep 3 00:45:25 2015 IgnacioUpdateIOORemeasured MC2 to MCL TF + Improved subtraction performance

Today, I remeasured the transfer function for MC2 to MCL in order to improve the subtraction performance for MCL and to quantify just how precisely it needs to be.

Here is the fit, and the measured coherence. Data is also attached here: TF.zip

 

OMG, I forgot to post the data and any residuals. LOL!

The transfer function was fitted using vectfit with a weighting based on coherence being greater than 0.95.

I then used the following filters to do FF on MCL online:

Here are the results:

Performance has definelty increased when compared to previous filters. The reason why I think we still have poor performance at 3 Hz, is 1) When I remeasured the transfer function, Eric and I were expecting to see a difference on its shape due to the whitening filters that were loaded a couple days ago. 2) Assuming the transfer function is correct, there is poor coherence at 3 Hz 3) The predicted IIR subtraction is worst at this frequency.

Attachment 1: TF.zip
  11562   Thu Sep 3 00:29:45 2015 ranaUpdateASCASS X not working

ran the ON script several times and it kept pulling it away from good alignment, even when TRX was > 0.9.

Also, for what reason was this model run at 16 kHz?? Makes no sense to me to have a low frequency servo system run so high. Only makes for more digital precision noise, more CPU time, etc. Of course, running it at 2k would mean having to think about all of the AA filtering needed to go up/down from 2k to 16k.

  11561   Thu Sep 3 00:14:09 2015 ranaConfigurationIOOIMC fast gain change for lock acq

The IMC often was making that scratchy noise when first catching lock and sometimes breaking. Thinking of the crappy crossover sit that EQ showed in his latest plots, I decided that it didn't make sense to acquire lock with an unstable PZT/EOM crossover, so I have changed mcdown to acquire with +13 dB Fast Gain and its much fast now and no longer makes that sound.

I also changed the caput command from 'caput -l' to 'caput -c -l' to see if the async 'wait for callback' feature will insure that the commands get sent. I witnessed the mcdown not actually writing all of its commands once or twice tonight. With the MC Boost left on its never going to lock.

mcdown has been committed to SVN. Please, if you have recently edit mcup and Autolock, commit them to the SVN or else I will delete them and do an svn up.

  11560   Wed Sep 2 23:50:00 2015 ranaUpdatePSLPMC LO dying

Let's order a pair of 35.5 MHz Wenzel for this guy and package like Rich has done for the WB low noise oscillators.

WE're only sending 6 dBm into it now and its using a 13 dBm mixer. Bad for PMC stability.

Also, if anyone has pix of the servo card, please add them to the DCC page for the PMC.

Attachment 1: PMCLO.png
PMCLO.png
  11559   Wed Sep 2 13:44:04 2015 ericqUpdateLSCIR beatnote confusion

From the AFW website about our product, the POBC-64-C-1-7-2-25dB:

port1 slow axis -> port3 slow axis 
port2 slow axis -> port3 fast axis

crying

  11558   Wed Sep 2 01:31:22 2015 ericqUpdateIOOIMC loop shapes

The promised historical comparisons follow. The crossover looks mostly the same as before. There is a new feature in the OLG at 50-60kHz; what could've changed about the EOM path in that time?

 

Attachment 1: imc_olg_trend.pdf
imc_olg_trend.pdf
Attachment 2: imc_xover_trend.pdf
imc_xover_trend.pdf
  11557   Tue Sep 1 17:38:58 2015 ericqUpdateLSCIR beatnote confusion
Quote:

Maybe we just don't understand the splitter/combiners.

After an email from Eric G, I think this is the case.

If you read the text at Thorlabs about Fiber-Based Polarization Beam Combiners/Splitters, it suggests that these things take input beams both aligned to their slow axes, and outputs one field along the slow, and one orthogonal to it on the fast axis. Which is exactly what we don't want for a beat. 

  11556   Tue Sep 1 17:07:06 2015 ericqUpdateLSCIR beatnote confusion

There has been some discussion here and there of using fiber coupled IR beats for ALS. A few weeks ago, and again today with Eric G, I poked around a bit with the fiber box Manasa set up for the FOL scheme. 

Somehow, the IR beatnote is ~1000 times smaller than expected, both with the Thorlabs fiber coupled PD and a fiber coupled NF 1611.

In essence, after the fiber combiner, there is on the order of hundreds of uW each of PSL and AUX X IR light. The output of the fiber from each source looks nice and gaussian. The DC output of the 1611 indicates that it is seeing the right level of light. The green beatnote exists with good SNR at twice the IR beat frequency, so we know that the IR beat isn't some junky modes beating. 

For the 1611, we would expect an RF signal of ~1mW*0.9A/W*700V/A -> .6V / +8dBm. Instead we see ~2mV / -40dBm.

Incidentally, there is some 20mV / -20dBm signal at ~400kHz, presumably from the green PDH modulation at ~200k. 

(The level out of the thorlabs PD is similarly tiny; it doesn't have a DC output though, so we don't know the DC power that the active surface really sees. Not that I expect it to be much different, but the NF just makes it easier to estimate.)

The only things that should be able to cause the beat to be smaller than expected from the power levels are mode matching and polarization matching. All the fibers are single mode, so mode matching should be effectively 100%. Maybe somthing fishy is happening with the polarizations, but they'd have to be really maliciously close to orthogonal to cause this level of mismatch. 

Maybe we just don't understand the splitter/combiners. Mysterious.

  11555   Tue Sep 1 11:56:56 2015 ericqUpdateIOOIMC loop shapes

I took some transfer functions of the IMC loop and crossover, being careful that the PC drive never exceeding 1V during the measurements. 

I then did some algebra to try and back out the individual loop paths, without having to make assumptions/approximations about the loop gain being high enough. This only really works in the region where both the open loop and crossover measurements have coherence. 

It seems to me that the PZT path has pretty low phase margin on its own, but maybe this is ok, since its never really meant to run solo. The EOM path shape is harder to understand.

 

The data I took, and code that made the above plot is attached. This afternoon, I'll post an update comparing the measured OLG and crossover to earlier measurements. 

Attachment 1: IMCshapes_Aug31_2015.pdf
IMCshapes_Aug31_2015.pdf
Attachment 2: IMC_Aug31_2015.zip
  11554   Tue Sep 1 10:36:06 2015 SteveUpdatePEMGurs swapped
Quote:

To help find out if Steve really melted the inside of our precious seismometer, lets hook it up using the handheld seismo wand and see if it produces volts when we shake the ground.

Also, please stop using names like GurA or Gur1 or GurSuzy. We have GurX and GurY because they are at those ends. Anything else is confusing.

I moved Gur A from ETMX to ETMY . Gur B at ETMY was disconnected and its cable  connected to Gur A

It seems that Gur A is alive. I will stop using A and B names after we stop swapping components.

Attachment 1: GurA@ETMY.png
GurA@ETMY.png
  11553   Tue Sep 1 10:26:24 2015 IgnacioUpdateIOOMore MCL Subtractions (Post FF)

Using the training data that was collected during the MISO MCL FF. I decided to look at more MCL subtractions but this time using the accelerometers as Rana suggested.

I first plotted the coherence between MCL and all six accelerometers and the T240-Z seismometer.

For 1 - 5 Hz, based on coherence, I decided to do SISO Wiener filtering with ACC2X and MISO Wiener filtering with ACC2X and ACC1Y. The offline subtractions were as follows (RMS plotted from 0.1 to 10 Hz):

The subtractions above look very much like what you would get offline when using the T240(X,Y) seismometeres during MISO Wiener filtering. But this data was taken with the MISO filters on. This sort of shows the performance deterioration when one does the online subtractions. This is not surprising since the online subtraction performance for the MISO filters, was not too great at 3 Hz. I showed this in some other ELOG but I show it again here for reference:


Anyways, foor 10 - 20 Hz, again based on coherence, I decided to do SISO Wiener filtering with ACC2Z and MISO Wiener filtering with ACC2Z and ACC1Z (RMS plotted from 10 to 20 Hz):

I will try out these subtractions online by today. I'm still debating wether the MISO subtractions shown here are worth the Vectfit shananigans. The SISO subtractions look good enough.

Attachment 4: mclxycoh.png
mclxycoh.png
  11552   Tue Sep 1 06:58:11 2015 IgnacioUpdateWienerFilteringMCL FF => WFS1 and WFS2 FF => ARMS FF

I took some training data during Sunday night/Monday morning while the MCL MISO FF was turned on. We wanted to see how much residual noise was left in the WFS1/WFS2 YAW and PITCH signals. 

The offline subtractions that can be achieved are:

For WFS1

 

For WFS2

I need to download data for these signals while the MCL FF is off in order to measure how much subtraction was achived indirectly with the MCL FF. In a previous elog:11472, I showed some offline subtractions for the WFS1 YAW and PITCH before any online FF was implemented either by me or Jessica. From the plots of that eLOG, one can clearly see that the YAW1 signal is clearly unchanged in the sense of how much seismic noise was mitigated indirectly torugh MCL. 

Koji has implemented the FF paths (thank you based Koji) necessary for these subtractions to be implemented. The thing to figure out now is where we want to actually actuate and to measure the corresponding transfer functions. I will try to have either Koji or Eric help me measure some of these transfer functions.

Finally, I looked at the ARMS and see what residual seismic noise can be subtracted

 

I'm not too concerned about noise in the arms as if the WFS subtractions turn out to be promising then I expect for some of the arms seismic noise to go down a bit further. We also don't need to measure an actuator transfer function for arm subtractions, give that its essentially flat at low frequencies, (less than 50 Hz).

 

  11551   Tue Sep 1 02:44:44 2015 KojiSummaryCDSc1oaf, c1mcs modified for the IMC angular FF

[Koji, Ignacio]

In order to allow us to work on the IMC angular FF, we made the signal paths from PEM to MC SUSs.
In fact, there already were the paths from c1pem to c1oaf. So, the new paths were made from c1oaf to c1mcs. (Attachment 1~3)

After some debugging those two models started running. The additional cost of the processing time is insignificant.
FB was restarted to accomodate the change.

Once the modification of the models was completed, the OAF screens were modified. It seemed that the Kissel button
for the output matrix haven't been updated for the PRM ASC implementation. This was fixed as the button was updated this time.
In addition, the button for the FM matrix was also made and pasted.

 

Attachment 1: c1oaf_screenshot1.png
c1oaf_screenshot1.png
Attachment 2: c1oaf_screenshot2.png
c1oaf_screenshot2.png
Attachment 3: c1mcs_screenshot.png
c1mcs_screenshot.png
Attachment 4: OAF_MEDM1.png
OAF_MEDM1.png
Attachment 5: OAF_MEDM2.png
OAF_MEDM2.png
  11550   Mon Aug 31 14:15:23 2015 IgnacioUpdateIOOMeasured the MC_F whitening poles/zeroes

I measured the 15 Hz zero and the 150 Hz pole for the whitening filter channels of the Generic Pentek board in the IOO rack. The table below gives these zero/pole pairs for each of the 8 channels of the board.

channel zero [Hz] pole [Hz] Chan
1 15.02 151.05 C1:ASC-POP_QPD_YAW
2 15.09 150.29 C1:ASC-POP_QPD_PIT
3 14.98 150.69 C1:ASC-POP_QPD_SUM
4 14.91 147.65 C1:ALS-TRX
5 15.03 151.19 C1:ALS-TRY
6 15.01 150.51 ---
7 14.95 150.50 C1:IOO-MC_L
8 15.03 150.93 C1:IOO-MC_F

Here is a plot of one of the measured transfer functions,

and the measured data is attached here: Data.zip


EQ: I've added the current channels going through this board. 

More importantly, I found that the jumpers on channel one (QPD X) were set to no whitening, in contrast to all other channels. Thus, the POP QPD YAW signals we've been using for who knows how long have been distorted by dewhitening. This has now been fixed. 

Hence, the current state of this board is that the first whitening stage is disabled for all channels and the second stage is engaged, with the above parameters. 

Attachment 1: Data.zip
  11549   Mon Aug 31 09:36:05 2015 IgnacioUpdateIOOMISO Wiener Filtering of MCL

MISO Wiener filters for MCL kept the mode cleaner locked for a good 8+ hours.

  11548   Mon Aug 31 07:49:11 2015 ericqUpdateIOOMC2 -> MCL Actuator TF
Quote:

I think what happened here is you forgot to undo the MC_F whitening filter which is the Generic Pentek Interface board next to the MC servo board. I suggest you guys measure this on Monday so you can correctly estimate the MC length noise. And then perhaps undo the whitening in the anti-whitening filter of this filter bank so that the signal which is recorded is in units of kHz.

This should allow your online subtraction filter to be more correct: roughly speaking, the phase shift below a pole or zero is going to be 45*(f/fp) deg. Since we expect there to be 2 zeros at 15 Hz, it would be 9 deg phase shift at 1.5 Hz and limit the subtraction to ~80%.

While it is true that the whitening filter was incorrectly handled, I don't think this should change the subtraction performance since the MC_L data used for the Wiener filter training was also taken without undoing the whitening filter.

ELOG V3.1.3-